I am having problems with solving by elimination...problem example:
-3a+2b=0
3a-4b=-1
Then I do not understand how this gives me an ordered pair.

Answer by jorel1380(3719)

Add the two together:
-3a+2b=0
3a-4b=-1
-2b=-1
b=1/2
a=1/3
Your ordered pair is (1/3,1/2)..